OpenStudy (ahvei):

i will help any and everyone but you have to place your question here and you HAVE TO MEDAL ME..........I WILL ANSWER AS FAST AS I CAN BUT YOUR GOING TO HAVE TO TAG ME IN IT LIKE @ahvei and i will come

OpenStudy (anonymous):

Find the absolute value and the opposite of the integer. 1534 A. The absolute value is 1534 and the opposite is –1534. B. The absolute value is –1534 and the opposite is –1534. C. The absolute value is –1534 and the opposite is 1534. D. The absolute value is 1534 and the opposite is 1534.

OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (mrnood):

I suggest anyone who feels like using this poster to help FIRST looks at previous ansers he/she has posted. I think you will find that there are more trustworthy people to help you in your education

OpenStudy (ccastudentgrade10):

PLEASE HELP ASAP write the equation in slope intercept form -9x + 10y = -9 @ahvei

OpenStudy (ahvei):

ok @MrNood you have 1 minute to get off of MY post of i will get a moderator and i will report you 100000 times

OpenStudy (ahvei):

A is the answer @aliyahbufford1239442

OpenStudy (ahvei):

was there a picture that came wit that @ccastudentgrade10

OpenStudy (mrnood):

@ccastudentgrade10 of course you do not need a 'picture' with that. slope-intercept form is a general form of the equation for a straight line and is always in the form y= mx +b where m and b are constants so oyu need to take your equation -9x + 10y = -9 and rearrange it so it looks like y= mx +b so first you need the y term on its own on one side of the equation so add 9x to each side of the equation. it will then look like 10y = ax + c so divide both sides by 10. You will then have your answer by maths - not by pictures.

geerky42 (geerky42):

Thank you, but you can just go to other questions posted here and answer them.

OpenStudy (sleepyjess):

@ahvei , direct answers are not allowed. If you're going to make a post like this, at least guide the asker to an answer :)

OpenStudy (jack1):

There are n sweets in a bag. Six of the sweets are blue. The rest of the sweets are purple. Hannah takes a sweet from the bag. She eats the sweet. Hannah then takes at random another sweet from the bag. She eats the sweet. The probability that Hannah eats two blue sweets is 1/3. how many blue and how many purple sweets in the bag?

OpenStudy (ahvei):

i always ask if they need help of just want the answer so bye

OpenStudy (ahvei):

is it a multiple chice

OpenStudy (sleepyjess):

Direct answers are not allowed

Nnesha (nnesha):

my question :=)

OpenStudy (ahvei):

ok bye dangggggggggg

Nnesha (nnesha):

\[\huge\rm \frac{ (2x)(4y^2)(5x^4)^0 (568)^3(0) }{ 0 }\]

OpenStudy (sleepyjess):

I forgot, I do need help on one question... Find a ⋅ b. a = <3, 4>, b = <7, 5>

OpenStudy (jack1):

ahvei plz hlp how many of each colour?

OpenStudy (ricktoraw):

help -15 over 2 3 +5 over 2 3

OpenStudy (alexandervonhumboldt2):

to find a*b @sleepyjess just substitute your values in the a*b: a*b=3.4*7.5=solve